Prove or disprove: If A, B, and C are nonempty sets, and A×B = A×C, then B = C.
1
Expert's answer
2019-09-23T07:12:11-0400

Let "A", "B", and "C" be nonempty sets such that "A\times B = A\times C". Since "A" is nonempty, there is "a\in A". Now we prove that "B = C". Let "x\in B". Then "\langle a, x\rangle\in A\times B" by the definition of set product. Since "A\times B = A\times C", "\langle a, x\rangle\in A\times C", and "x\in C" again by the definition of set product. Therefore, "B" is included in "C". By a symmetric proof, "C" is included in "B". Therefore, "B = C". (Actually, the assumption that "B" and "C" are nonempty is not needed.)